Using the married filing jointly status and their income and expense​ statement, calculate the

2017

tax liability for Shameka and Curtis Williams.​ First, use the standard​ deduction, and then use the following itemized​ deductions:

Income

Expenses

Earned income

​$49,000.00

Home mortgage interest

​$7,800.00

Interest income

1,700.00

Real estate and state income taxes

3,800.00

Miscellaneous deductions

750.00

Explain to the Williams which method they should use and why.

Shameka and​ Curtis' total gross income for the

2017

tax year is

​$nothing.

​(Round to the nearest​ cent.)Assuming Shameka and Curtis are filing​ jointly, their exemption amount for the

2017

tax year is

​$nothing.

​(Round to the nearest​ cent.)If Shameka and Curtis use the standard​ deduction, their standard deduction amount for the

2017

tax year is

​$nothing.

​(Round to the nearest​ dollar.)  ​Note: Miscellaneous​ deductions, unlike charitable contributions are subject to a deduction​ threshold.) Click the following link for a

standard

deduction table.

LOADING...

Their taxable income for the

2017

tax year is

​$nothing.

​(Round to the nearest​ cent.)If Shameka and Curtis use the standard​ deduction, their total income tax due for the

2017

tax year is

​$nothing.

​(Round to the nearest​ cent.) Click the following link for the tax rates and brackets table.

LOADING...

If Shameka and Curtis itemize their​ deductions, their itemized deduction amount for the

2017

tax year is

​$nothing.

​(Round to the nearest​ cent.)If Shameka and Curtis itemize their​ deductions, their taxable income for the

2017

tax year is

​$nothing.

​(Round to the nearest​ cent.)If Shameka and Curtis itemize their​ deductions, their total income tax due for the

2017

tax year is

​$nothing.

​(Round to the nearest​ cent.)  Click the following link for the tax rates and brackets table.

LOADING...

Which method should they​ use?  ​(Select the best answer​ below.)

A.

Itemized deduction.

B.

Standard deduction.
Answer #1

Amounts are in $

1. Gross Total Income

Earned Income = 49,000

Interest Income = 1,700

Gross Total Income = 50,700

2. Exemption amount for 2017 tax year is $4,050 per individual

So total exemption amount = $8,100

3. Standard deduction maximum limit for married tax payers filing jointly for 2017 tax year is $12,700

4. Taxable income if they opt for standard deduction

Taxable income = Gross Income - Personal exemption - Standard deduction

= $50,700 - $8,100 - $12,700

= $29,900

5. Tax on the above taxable income

= 1,865 + 15% (29,900-18,650)

= 1,865 + 1,687.5

= 3,552.5

6. Itemised deduction amount

= Mortgage interest + Real-estate taxes

= 7,800 + 3,800

= 11,600

7. Taxable income if we itemize deduction

Taxable Income = Gross Income - Personal exemption - Itemized deduction amount

= $50,700 - $8,100 - $11,600

= $31,000

8. Tax on above income

= 1,865 + 15% x (31,000 - 18,650)

= 1,865 + 1,852.5

= $3,717.5

9. As the Tax is lower ($3,552.5) when we take standard deduction when compared to itemizing our deduction ($3,717.5), we opt to take standard deduction

Note : There is a miscellaneous deduction of $750, as there is no enough details on such amount we have taken that it as ineligible for Itemized deduction